if y= 1+(a/x-a)+(bx/(x-a)(x-b))+(cx^2/(x-a)(x-b)(x-c))
than show that
dy/dx=y/x[(a/a-x)+(b/b-x)+(c/c-x)].............

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear pradeep,

[a/(x-a)]+1=x/(x-a)
[bx/(x-b)(x-c)]+x/(x-a)=x/(x-a)[b/(x-b)+1]=x2/(x-a)(x-b)
So y={cx^2/ (x-a) (x-b) (x-c)} +x2/(x-a)(x-b)=x2/((x-a)(x-c))[c/(x-c)+1]
y=x3/(x-a)(x-b)(x-c)
logy=3logx-log(x-a)-log(x-b)-log(x-c)
y1/y=3/x+1/(a-x)+1/(b-x)+1/(c-x)
y1/y=[1/x-1/(x-a)]+[1/x-1/x-b]+[1/x-1/x-c]=1/x ( a/a-x + b/ b-x+ c/c-x)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

y = 1 + a/(x-a) +bx/(x-a)(x-b) + cx2/(x-a)(x-b)(x-c)

breaking into different  functions

 

y = 1 +    f(x)1  +       f(x)2       +        f(x)3       

 

f(x)1=a/(x-a)    ,        f(x)2 =bx/(x-a)(x-b)   &   f(x)3 =cx2 /(x-a)(x-b)(x-c)

now

          dy/dx= d/dx(1) + d/dx(f(x)1) +d/dx(f(x)2) +d/dx(f(x)3) ....................................1

now separatly differentiatiating f(x)1 , f(x)2  , f(x)3  by taking log

  (taking log in both sides)

          d/dx[log(f(x)1)]  =  d/dx(loga/x-a)

                d/dx(f(x)1/f(x)1 = d/dx(loga - log(x-a) )                 

 

                     d/dx(f(x)1)=(-1/x-a)(f(x)1

                       d/dx(f(x)1) =    -a/(x-a)2

now separatly differentiate each of these and put in eq 1 ,u will get the required result